OpenStudy (hayhay12345):

what is an equation of the line in slope intercept form m=3 and the y intercept is (0,4)

OpenStudy (hayhay12345):

y=4x+3 y=-3x+4 y=4x-3 y=3x+4

OpenStudy (hayhay12345):

i think its the last one

OpenStudy (candyme):

The slope intercept form is y=mx+b. Since m is the slope, we can fill in y=3x+b We also know that b is the y part of the y intercept. SO the full equation is y=3x+4

OpenStudy (candyme):

Great! You had it!

OpenStudy (hayhay12345):

i got it right?! sweet

OpenStudy (hayhay12345):

can u help me with another?

OpenStudy (candyme):

Sure!

OpenStudy (hayhay12345):

im going to screen shot it

OpenStudy (candyme):

This one will be much easier if you get the y all by itself on the left. So start by adding the x to both sides. -12y-5x+5x=5x+11 -12y=5x+11 Now you want to divide by 12\[y=-\frac{ 5 }{ 12 }x-\frac{ 11 }{ 12 }\]

OpenStudy (candyme):

So the slope is the number in front of the x \[-\frac{ 5 }{ 12 }\] And the y intercept is the number that has no variable\[-\frac{ 11 }{ 12 }\]

OpenStudy (hayhay12345):

so it is the third one right?

OpenStudy (candyme):

That's right!

OpenStudy (hayhay12345):

thanks!! do u miind checking a few more for me i am pretty sure they are right i just want to make sure.

OpenStudy (candyme):

Sure

OpenStudy (candyme):

Ok, on the first one the picture cut off that last piece of the equation. Based on your slope the answer is either b or c, but you need that number on the right side of the equals sign to finish it

OpenStudy (candyme):

The second one you posted you are correct

OpenStudy (hayhay12345):

for the first on it was x+2y=4

OpenStudy (candyme):

For this one there is no slope, so there won't be a number with the x. Vertical lines are always written as x=__. So this graph is x=2

OpenStudy (candyme):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @hayhay12345 for the first on it was x+2y=4 \(\color{#0cbb34}{\text{End of Quote}}\) Ok, so when you divide by 2 you will have \[y=-\frac{ 1 }{ 2 }x+2\]

OpenStudy (candyme):

So on it your graph would be c because it hits the y axis at 2

OpenStudy (hayhay12345):

ohhhh ok thank you!! can u take a look at these two?

OpenStudy (candyme):

For the vertical line above, you always write the equation as x=__. Since it crosses the x axis at (2,0) you write x=2

OpenStudy (hayhay12345):

i put the wrong one sorry

OpenStudy (hayhay12345):

its this one

OpenStudy (candyme):

Both of those are correct :)

OpenStudy (hayhay12345):

sweet i hate to ask again but could u check like 4-5 more for me plz?

OpenStudy (candyme):

I can do that :)

OpenStudy (hayhay12345):

thats for one of them

OpenStudy (hayhay12345):

i had b for 14 btw

OpenStudy (candyme):

For that first one it is going to be d because it makes it skinny and it isn't negative so it won't flip upside down

OpenStudy (candyme):

For 14 you are correct

OpenStudy (candyme):

Also letter b for the one after that, because the > means dashed line and shaded above

OpenStudy (hayhay12345):

thats all other than what i need to finish working out

OpenStudy (hayhay12345):

i just need those last one i posted checked before i finish the complete work

OpenStudy (candyme):

Sorry I froze up there for a few minutes. I'll look at these for you real quick

OpenStudy (hayhay12345):

ok and its all cool i was working out some others ones

OpenStudy (candyme):

The first one should actually be d because you want the shading to be on the bottom area.

OpenStudy (hayhay12345):

ohh ok

OpenStudy (candyme):

On the absolute value graph, you need to move that 4 over to the other side. That will make it y=|x+2|-4 so it is actually c

OpenStudy (candyme):

The one about DVDs was correct

OpenStudy (hayhay12345):

@candyme can you check these for me?

OpenStudy (hayhay12345):